OpenStudy (anonymous):

What is (-m)^-3 n If m=2 and n=-24 What is -a^-2 if a=-5 ^ Means exponent also what simplifies to a negative # -4^-4 ,(-4)^-4,4^-4,1/4^-4

OpenStudy (anonymous):

If any one answers correctly I will give them a medal !

OpenStudy (anonymous):

-m)^(-3)* n (-2)^(-3) *(-24) a^(-b)=1/a^(b) (-2)^(-3)=1/(-2)^3 1/(-2)^3 * -24 -24/ -8= 3 So 3.